OpenStudy (kkutie7):

find the value of k for which \[y=x^{2}+k\] is a solution to the differential equation: \[2y-xy'=28\] my work below

OpenStudy (kkutie7):

\[2y-x'y=28\rightarrow x\frac{dy}{dx}=2y-28\rightarrow \] \[\frac{1}{2}\int\limits_{}^{}\frac{ dy }{ y-14 }=\int\limits\frac{dx}{x}\rightarrow \] \[y=\frac{x+c+14}{e^{\frac{1}{2}}}\] this is where I get confused.

OpenStudy (superdavesuper):

err it will be easier if u go the OTHER way....by differeniating y....

OpenStudy (kkutie7):

uh ok...

OpenStudy (superdavesuper):

y'=2x n put it back in the diff eqn...

OpenStudy (superdavesuper):

oh it should be y'=2x AND y=x^2 + k into the DE

OpenStudy (kkutie7):

so the answer is 14

OpenStudy (superdavesuper):

@Kkutie7 correct - good job! :)